OpenStudy (anonymous):

Solve. Check for extraneous solutions. sqrt2x+6 - sqrtx-1=2

OpenStudy (anonymous):

\[\sqrt{2x+6}-\sqrt{x-1}=2\] @Hero

OpenStudy (anonymous):

Alright, any idea how to start?

OpenStudy (anonymous):

I have no clue, that's why I need it explained in detail.

OpenStudy (anonymous):

Well first lets start off by getting rid of the squareroots on the left side, you can do that by squaring both sides.

OpenStudy (anonymous):

\[(\sqrt{2x+6}-\sqrt{x-1})^2 = 2^2\] What do you get?

OpenStudy (anonymous):

I'm not positive, I know it has to be x=

OpenStudy (the_fizicx99):

Whale, obviously, you're solving for the only variable C: What's \(\ (\sqrt{2x+6}-\sqrt{x-1})^2 = 2^2 \) ?

OpenStudy (anonymous):

I don't know, I've tried 2, 0, 3, 5, and 10 in place of x

OpenStudy (the_fizicx99):

Square both sides is a good start

OpenStudy (anonymous):

i don't understand why you're plugging in numbers :o, we have to solve for x.

OpenStudy (the_fizicx99):

*Cough* You want to get rid of the square root, so you do the inverse, which is squaring it. Squaring both sides, you get? >.>

OpenStudy (anonymous):

I'm so confused

hero (hero):

\(\sqrt{2x+6}-\sqrt{x-1}=2\) First add \(\sqrt{x - 1}\) to both sides

hero (hero):

@ChasingMaggie, are you still here?

OpenStudy (anonymous):

Yes @Hero

hero (hero):

Did you add \(\sqrt{x - 1}\) to both sides?

OpenStudy (anonymous):

I really don't think I'm following. Would the equation then be \[\sqrt{2x+6} - \sqrt{x-1}+\sqrt{x-1}=2+\sqrt{x-1}\]

hero (hero):

Yes and obviously \(-\sqrt{x - 1} + \sqrt{x - 1} = 0\) so you have \(\sqrt{2x + 6} = 2 + \sqrt{x - 1}\)

OpenStudy (anonymous):

And the first part of the equation would then turn into \[\sqrt{2x+6}=\]

hero (hero):

Next square both sides: \((\sqrt{2x + 6})^2 = (2 + \sqrt{x - 1})^2\)

hero (hero):

What will the left side simplify to?

OpenStudy (anonymous):

Would it be \[\sqrt{4x+12}\]

hero (hero):

Actually, in case you no one told you, when you apply a square to a square root, they will cancel since square and square root are inverses of each other. In this case, for example, \(\sqrt{x^2} = x\)

hero (hero):

Likewise \((\sqrt{2x + 6})^2 = 2x + 6\)

hero (hero):

So the left side of the equation simplifies to \(2x + 6 = (2 + \sqrt{x - 1})^2\)

hero (hero):

Meanwhile on the right side, we must expand: \(2x + 6 = (2 + \sqrt{x - 1})(2 + \sqrt{x - 1})\)

OpenStudy (anonymous):

What's next?

hero (hero):

On the right side, continue expanding. Your class might call it FOIL. Here we just call it multiplying binomials or finding the product of two binomials.

OpenStudy (anonymous):

I never learned that. I tested out of my other math class into this one last week, and I'm expected to know all this.

hero (hero):

So you never learned how to multiply something like (x + 2)(x + 3)?

OpenStudy (anonymous):

No, I know how to do that.

hero (hero):

Are you familiar with the distributive property?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

would it be \[2x+6=(\sqrt{x}+1)^{2}\]

OpenStudy (anonymous):

never mind, that's not right

OpenStudy (anonymous):

@hero can you just help me through the rest?

hero (hero):

It's kind of difficult to help you if you're not familiar with distributive property.

hero (hero):

What about 2(x + 5) do you know how to expand that?

hero (hero):

By the way distributive property is a(b + c) = ab + ac